factoring a quadratic with leading coefficient greater than one
2y^2+15y+28

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(75887)

factoring a quadratic with leading coefficient greater than one
2y^2+15y+28
ac=2*28=56
b=15
Think of two numbers whose product is 56 and whose sum is 15
The numbers are 7 and 8
Replace 15y by 7y+8y to get:
2y^2+7y+8y+28
Factor to get:
y(2y+7)+4(2y+7)
Factor again to get:
(2y+7)(y+4)
=============
Cheers,
Stan H.

Answer by checkley71(8403)

2Y^2+15Y+28
FACTORS OF 2=2*1 & FACTORS OF 28=(28*1),(14*2),(7*4)
NOW FIND THE SUM OF THE PRODUCTS OF THESE FACTORS THAT=15.
2*28+1*1=56+1=57 NOT A SOLUTION
2*4+1*7=8+7=15 LOOKS LIKE A SOLUTION.
(2X+7)(X+4)